OpenStudy (timothywhitaker):

does anyone know how to put y=3x+4 into standard form

OpenStudy (gorv):

ax+by=c is standard form

OpenStudy (gorv):

subtract 3x from both side

OpenStudy (gorv):

y-3x=4

OpenStudy (gorv):

-3x+y=4

OpenStudy (gorv):

multiply by -1 on each side ..what u got ??

OpenStudy (anonymous):

The standard form is -3x+y=4